# A Study on the Calibration of In-context Learning

### 07 Dec 2023 v1

@@ Abstract
Modern auto-regressive language models are trained to minimize log loss on broad data by predicting the next token so they are expected to get calibrated answers when framing a problem as next-token prediction task. We study this for in-context learning (ICL), a widely used way to adapt frozen large language models (LLMs) via crafting prompts, and investigate the trade-offs between performance and calibration on a wide range of natural language understanding and reasoning tasks. We conduct extensive experiments to show that such trade-offs may get worse as we increase model size, incorporate more ICL examples, and fine-tune models using instruction, dialog, or reinforcement learning from human feedback (RLHF) on carefully curated datasets. Furthermore, we find that common recalibration techniques that are widely effective such as temperature scaling provide limited gains in calibration errors, suggesting that new methods may be required for settings where models are expected to be reliable.
@@ 1 Introduction
In-context learning has been expected to learn models by gradient descent in their forward pass (Von Oswald et al., 2023), which might hopefully yield calibrated predictions (Błasiok et al., 2023) if the models are getting close to local optimality with respect to test loss through meta-optimization. However, the fact that choosing ICL samples from the validation set does not naturally lead to calibrated predictions shows that ICL learns in a fairly different way than SGD. We design controlled experiments to illustrate task learning properties of ICL, showing that when examples in the prompt demonstrate consistent task properties, the learning performance, and calibration would be improved.
@@ 2 Related Work
Uncertainty quantification in NLP. Uncertainty quantification in NLP, which often adopts the Bayesian principle to sophisticated methods tailored for neural networks, aims to enhance the reliability of model predictions. This may involve non-trivial designs as directly interpreting language model predictions via probabilities (Kadavath et al., 2022) and linguistic expressions (Lin et al., 2022; Mielke et al., 2022; Zhou et al., 2023) may inadvertently lead to over-reliance on the model’s uncertainties (Si et al., 2023), thus complicating the establishment of trustworthy common ground between humans and models (Buçinca et al., 2021). Notable recent advancements include employing model confidence as a critical factor in various applications like dialogue generation (Mielke et al., 2022), cascading prediction (Schuster et al., 2021), open-domain QA (Fisch et al., 2020; Angelopoulos et al., 2022), summarization (Laban et al., 2022), language modeling (Schuster et al., 2022), image captioning (Petryk et al., 2023).
@@ 3 Background
Refer to caption (a) 7B Refer to caption (b) 13B Refer to caption (c) 30B
Figure 2: Reliability plots of LLaMA models.
@@
Confidence and feature norm. We record the maximum probability of the answer token as its confidence $\textrm{Conf}=\mathcal{P}_{\theta}(y_{n}|w_{<n})$ and the feature norm $z_{n}$ as the hidden states of the answer token from the output of the last layer of the model.
@@ 4.1 Experimental Settings
Models. We study decoder-only autoregressive LMs involving GPT-2, LLaMA (Touvron et al., 2023a), and their variants fine-tuned with instruction, dialog, or RLHF like Alpaca (Dubois et al., 2023), Vicuna (Zheng et al., 2023), and LLaMA2-Chat (Touvron et al., 2023b).
@@
In-context learning settings. We prompt the model via sampling $k$ examples from the training set for each test example in the $k$-shot setting. Each experiment is repeated 10 times to reduce variance and we report the mean results. We use $M=10$ bins for calculating calibration errors.
@@ 4.2 Numerical Results
The performance of LLaMA. We seek to characterize the calibration-accuracy trade-off in both simple and realistic settings (Tab. 1). We record the performance and calibration errors in both miscalibrated and recalibrated settings. Moreover, we take a close look at the prompting approaches that explicitly include explanations in reasoning tasks such as scratchpad (Nye et al., 2021) or chain-of-thought (Wei et al., 2022), showing that the calibration degrades after generating a long context for reasoning and explaining the final answer.
@@
  1. 1.
    We learn one temperature for each $n$-shot ICL, i.e., we learn different temperatures for different shot numbers in ICL;
  2. 2.
  3. 3.
Figure 3: Accuracy and calibration errors of LLaMA and its finetuned variants.
Looking into Fig. 6, none of the above strategies achieves satisfactory calibration performance, which is in contrast to the well-studied supervised learning setting where scaling the confidence scores (via temperature scaling) can effectively reduce calibration errors (Guo et al., 2017). The fact that applying a post-processing calibration method, such as temperature scaling, as used in most previous work, cannot directly resolve the miscalibration issue suggests that ICL might have substantially different properties compared to making predictions via classical supervised learning models, thus future investigations are needed to address such miscalibration issues.
The effect of finetuning. We show that vicuna and alpaca are both more accurate but less calibrated than their LLaMA counterpart backbones, the margin is especially large for reasoning tasks and vicuna. Thus we compare those models’ accuracy and ECE in Fig. 3, showing that finetuning might significantly degrade calibration, corroborating the evidence shown in (OpenAI, 2023), albeit it can improve the reasoning accuracy dramatically. Our results provide evidence that though fine-tuned on carefully curated datasets can greatly improve question-answering performance, especially for hard tasks like reasoning problems, attention may need to be paid when assessing the calibration of those models’ predictions.
The accuracy is high for 0-shot ICL but has not increased much as we include more in-context examples. We also note that the pattern of zero-shot performance is totally different for two fine-tuned models, i.e. vicuna, and alpaca.
The effect of prompt repetition. In our study investigating the impact of various prompt strategies, we employ three distinct approaches: Repeat-context: In this strategy, we construct the prompt as $w_{0}=[x_{1},x_{1},...,x_{1},y_{1}]$, where we repetitively include only the context $x_{1}$ a total of n times, excluding the label $y_{1}$ from repetition. Repeat-prompt: Here, we shape the prompt as $w_{0}=[x_{1},y_{1},...,x_{1},y_{1}]$, repeating both the context $x_{1}$ and the label $y_{1}$ n times within the prompt. Normal: In this strategy, we construct the prompt as $w_{0}=[x_{1},y_{1},x_{2},y_{2},...,x_{n-1},y_{n-1},x_{n},y_{n}]$, where distinct context-label pairs are systematically chosen to form the prompt. The results presented in Table 3 unveil essential insights: (1) The inclusion of labels within the prompt contributes to a reduction in uncertainty and facilitates more effective reasoning. Conversely, merely repeating the context without incorporating labels fails to yield improved performance. (2) Notably, the diversity inherent in the prompt’s construction significantly impacts performance, particularly concerning larger language models.
Table 3: Accuracy and Calibration of LLaMA-7B model and GPT-2 with different prompt repetition strategies.
Model Strategy 1-shot 4-shot 8-shot AVG ACC AVG ECE
Acc ECE Acc ECE Acc ECE
Repeat-Prompt 78.200 0.281 89.470 0.395 90.330 0.403 86.000 0.360
Normal 78.200 0.281 95.200 0.452 96.130 0.462 89.843 0.398
GPT-2 Repeat-context 65.970 0.287 55.300 0.230 54.330 0.230 58.533 0.249
Repeat-Prompt 65.970 0.287 62.500 0.147 57.970 0.176 62.147 0.204
Normal 65.970 0.287 63.030 0.212 56.270 0.220 61.757 0.240
@@ 4.3 Qualitative Results
We showcase that in SST-2 (4-shot), showing that both ACC and ECE of LLaMA increase as the model size increases (Fig. 2). We can observe that confidence scores tend to concentrate on values above 0.8 as we enlarge model sizes.
@@ 4.4 Ablation Studies
Ablation with model sizes. As we enlarge the sizes of models, they will become more confident and accurate (Fig. 2). As a result, the entropy decreases and ECE increases, showing that token-level calibration might have an inverse scaling relationship with model sizes.
@@
A closer look at the hidden state and confidence score. To better understand the miscalibration issue of ICL, we conduct fined-grained experiments to take a closer look at ICL properties: we measure the norm of the representation vectors [^2] The representation vector refers to the intermediate output before the linear prediction layer. for different number of shots in ICL, to better understand how the representation vectors are changing when increasing the number of shots in ICL. Meanwhile, we also measure the confidence and entropy of the prediction for $y_{n}$, and the results are summarized in Table 2. When switching from 0-shot to 1-shot, all three measurements (representation norm, entropy, and confidecent) drastically change. Meanwhile, more ICL samples lead to smaller entropy and higher confidence in most cases.
Confidence and wrongly classified reasoning examples. To take a closer look at the failure modes of LMs, we randomly sample 100 reasoning examples of LLaMA and plot the distribution of wrongly predicted samples and the confidence scores via thresholding. Similar to previous observations, as model sizes and the number of ICL examples scale up, LMs would generate more confident samples (Fig. 4 (c, d)). Note, that we observe "inverse scaling" behaviors where models with larger sizes are more error-prone and tend to generate more confidently wrong samples (Fig, 5).
Examples of hallucinated explanations for highly confident predictions. Next, we showcase in Table 4 that models generate both wrong explanations and incorrect predictions with high confidence. We also observe that most of the wrong predictions are highly confident. We manually examine the correctness of explanations on commonsense QA, and found its high correlations with predicted answer accuracy, which is the opposite of token-level explainability that tends to get worse when the accuracy improves.
@@ 5 Discussion and Concluding Remarks
In our investigation of the token-level calibration of in-context learning in contemporary Large Language Models (LLMs), we have delineated the intricate balance between ICL performance and calibration. Our findings underscore the importance of being circumspect in model deployment, as maximizing ICL performance does not invariably translate to improved calibration. As LMs continue to evolve and gain more capabilities such as having long enough context windows that can include the whole training set as in-context examples for some downstream tasks, our result can be pedagogical when users would like to examine their uncertainty through prediction probabilities. Moreover, the work suggests the following future directions:
@@ Acknowledgment
We thank Jean-Stanislas Denain for valuable feedback. Kempner Institute computing resources enabled this work.

### 11 Dec 2023 v2

@@ Abstract
Modern auto-regressive language models are trained to minimize log loss on broad data by predicting the next token so they are expected to get calibrated answers in next-token prediction tasks. We study this for in-context learning (ICL), a widely used way to adapt frozen large language models (LLMs) via crafting prompts, and investigate the trade-offs between performance and calibration on a wide range of natural language understanding and reasoning tasks. We conduct extensive experiments to show that such trade-offs may get worse as we increase model size, incorporate more ICL examples, and fine-tune models using instruction, dialog, or reinforcement learning from human feedback (RLHF) on carefully curated datasets. Furthermore, we find that common recalibration techniques that are widely effective such as temperature scaling provide limited gains in calibration errors, suggesting that new methods may be required for settings where models are expected to be reliable.
@@ 1 Introduction
In-context learning has been hypothesized as meta-learning by gradient descent in their forward pass (Von Oswald et al., 2023), which might hopefully yield calibrated predictions (Błasiok et al., 2023) if the models are getting close to local optimality with respect to test loss through meta-optimization. However, the fact that choosing ICL samples from the validation set does not naturally lead to calibrated predictions shows that ICL learns in a fairly different way than SGD. We design controlled experiments to illustrate task learning properties of ICL, showing that when examples in the prompt demonstrate consistent task properties, the learning performance would be improved while the accuracy-calibration trade-off still persists.
@@ 2 Related Work
Uncertainty quantification in NLP. Uncertainty quantification in NLP, which often adopts the Bayesian principle to sophisticated methods tailored for neural networks, aims to enhance the reliability of model predictions. This may involve non-trivial designs as directly interpreting language model predictions via probabilities (Kadavath et al., 2022) and linguistic expressions (Lin et al., 2022; Mielke et al., 2022; Zhou et al., 2023b) may inadvertently lead to over-reliance on the model’s uncertainties (Si et al., 2023), thus complicating the establishment of trustworthy common ground between humans and models (Buçinca et al., 2021). Notable recent advancements include employing model confidence as a critical factor in various applications like dialogue generation (Mielke et al., 2022), cascading prediction (Schuster et al., 2021), open-domain QA (Fisch et al., 2020; Angelopoulos et al., 2022), summarization (Laban et al., 2022), language modeling (Schuster et al., 2022), image captioning (Petryk et al., 2023).
@@
In-context Learning. Large models such as GPT-3 (Brown et al., 2020) have demonstrated the potential of in-context learning, a method where the model infers the task at hand from the context provided in the input, without requiring explicit retraining or fine-tuning for each new task. Some recent works attempt to understand ICL through meta-learning (Von Oswald et al., 2023), Bayesian inference (Xie et al., 2021), mechanistic interpretability (Olsson et al., 2022), algorithm selection (Bai et al., 2023), synthetic data and simple function classes (Garg et al., 2022; Akyürek et al., 2022). Notably, unlike previous works (Zhao et al., 2021; Han et al., 2023; Fei et al., 2023; Zhou et al., 2023a) that focus on improving task accuracy using the same “calibration” terminology, we study the uncertainty of ICL and measure its trade-off with accuracy.
@@ 3 Background
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 2: Reliability plots and confidence histograms of LLaMA models on 4-shot learning tasks. Results of different sizes 7B (left), 13B (middle), and 30B (right) are plotted.
@@
Confidence and feature norm. We record the maximum probability of the answer token as its confidence $\textrm{Conf}=\mathcal{P}_{\theta}(y_{n}|w_{<n})$ and the feature norm $z_{n}$ as the intermediate hidden state before the linear prediction layer.
@@ 4.1 Experimental Settings
Models. We study decoder-only autoregressive LMs involving LLaMA (Touvron et al., 2023a), and its variants fine-tuned with instruction, dialog, or RLHF like Alpaca (Dubois et al., 2023), Vicuna (Zheng et al., 2023), and LLaMA2-Chat (Touvron et al., 2023b).
@@
In-context learning settings. For $k$-shot learning, we prompt the model via sampling $k$ examples from the training set for each test example. Each experiment is repeated 10 times to reduce variance and we report the mean results. We use $M=10$ bins for calculating calibration errors.
@@ 4.2 Numerical Results
The performance of LLaMA. We record the performance and calibration errors for $k$-shot learning ($k=0,1,4,8$), characterizing the calibration-accuracy trade-off in both classic and realistic settings (Tab. 1). Our findings are two-fold: as more in-context examples are included, we observe a concurrent rise in both accuracy and calibration error across most situations, indicating a diminishing reliability of token-level probability in high performance ICL regimes; moreover, when $k=0$ increases to $k=1$, there is a marked boost in both accuracy and calibration error, demonstrating the importance of in-context examples and the significance of the established trade-off. In particular, for reasoning tasks, we explore prompting approaches that explicitly include explanations in reasoning tasks, i.e. scratchpad (Nye et al., 2021) or chain-of-thought (Wei et al., 2022), showing that calibration degrades after generating a long context for reasoning and explaining the final answer.
Table 2: Accuracy and calibration of LLaMA-7B model with different prompt repetition strategies.
Model Strategy 1-shot 4-shot 8-shot AVG ACC AVG ECE
Acc ECE Acc ECE Acc ECE
Repeat-prompt 78.200 0.281 89.470 0.395 90.330 0.403 86.000 0.360
Normal 78.200 0.281 95.200 0.452 96.130 0.462 89.843 0.398
  1. 1.
    We learn one temperature for each $k$-shot ICL, i.e., we learn different temperatures for different shot numbers in ICL;
  2. 2.
  3. 3.
Figure 3: Accuracy and calibration errors of LLaMA and its fine-tuned variants.
Looking into Fig. 6, none of the above strategies achieves satisfactory calibration performance, which is in contrast to the well-studied supervised learning setting where scaling the confidence scores (via temperature scaling) can effectively reduce calibration errors (Guo et al., 2017). The fact that applying a post-processing calibration method, such as temperature scaling, cannot directly resolve the miscalibration issue suggests that ICL might have substantially different properties compared to making predictions via classical supervised learning models.
The effect of fine-tuning. We show that vicuna, alpaca and LLaMA2-Chat are all more accurate but less calibrated than their LLaMA counterpart backbones (Fig. 3), the margin is especially large for reasoning tasks and vicuna. Our finding indicates that fine-tuning might significantly degrade calibration, corroborating the evidence reported in GPT-4 (OpenAI, 2023), albeit it can improve the reasoning accuracy dramatically. Our results provide evidence that though fine-tuned on carefully curated datasets can greatly improve question-answering performance, especially for hard tasks like reasoning problems, attention may need to be paid when assessing the calibration of those models’ predictions. We also note that the pattern of zero-shot performance is fairly different for two fine-tuned models, i.e. vicuna, and alpaca.
The effect of prompt formats. In our study, we explore the effects of different prompt strategies using three distinct methods. First, the Repeat-context approach involves constructing prompts as $w_{0}=[x_{1},x_{1},...,x_{1},y_{1}]$, where the context $x_{1}$ is repeated n times, but the label $y_{1}$ is not included in the repetition. Next, the Repeat-prompt strategy shapes the prompt as $w_{0}=[x_{1},y_{1},...,x_{1},y_{1}]$, where both the context $x_{1}$ and the label $y_{1}$ are repeated n times. Finally, the Normal involves constructing the prompt as $w_{0}=[x_{1},y_{1},x_{2},y_{2},...,x_{n-1},y_{n-1},x_{n},y_{n}]$, systematically incorporating distinct context-label pairs.
The findings, as detailed in Tab. 2, reveal certain insights: firstly, integrating labels within prompts significantly decreases uncertainty and enhances reasoning performance. In contrast, simply repeating the context without labels does not lead to better outcomes. Secondly, the diversity of ICL examples in prompt construction notably affects performance, especially for larger language models. Those observations corroborate that ICL is performant when the number of ICL examples is large and they demonstrate consistent task properties. Importantly, the trade-off persists for such controlled scenarios, i.e. improving learning performance in either way would suffer from poorer calibration.
@@ 4.3 Qualitative Results
We showcase that in SST-2 and Open Book QA (4-shot), both ACC and ECE of LLaMA increase as the model size increases (Fig. 2). Moreover, it’s observed that confidence scores tend to concentrate on high values as we enlarge the model size. In particular, the confidence of almost all predictions are above 0.8 for the 30B model.
@@ 4.4 Ablation Studies
Ablation with model sizes. As we enlarge the size of models, they become more confident and accurate (Fig. 2). As a result, the entropy decreases and ECE increases, showing that token-level calibration might have an inverse scaling relationship with model sizes.
@@
Examine the hidden state and confidence score. To better understand the miscalibration issue of ICL, we conduct fined-grained experiments to examine ICL properties: we measure the norm of the representation vectors [^2] The representation vector refers to the intermediate output before the linear prediction layer. for different number of shots in ICL. Meanwhile, we also measure the confidence and entropy of the prediction for $y_{n}$, and the results are summarized in Tab. 3. When switching from 0-shot to 1-shot, all three measurements (representation norm, entropy, and confidence) drastically change; on the other hand, when $k$ increases from $1\rightarrow 4\rightarrow 8$, the change of measures would become smoother. Our discovery shows that adding in-context example can substantially impact model behaviors while model behaves relatively similar for various shots once the task is specified ($k\neq 0)$. Meanwhile, more ICL samples lead to smaller entropy and higher confidence in most cases.
Confidence and wrongly classified reasoning examples. To inspect the failure modes of LMs, we randomly sample 100 reasoning examples of LLaMA and plot the distribution of wrongly predicted samples and the confidence scores via thresholding. Similar to previous observations, as model sizes and the number of ICL examples scale up, LMs would generate more confident samples (Fig. 4 (c, d)). We observe “inverse scaling” behaviors where models with larger sizes are more error-prone and tend to generate more confidently wrong samples (Fig. 5).
Examples of hallucinated explanations for highly confident predictions. Next, we showcase in Tab. 4 that models generate both wrong explanations and incorrect predictions with high confidence. We also observe that most of the wrong predictions are highly confident, thus we manually examine the correctness of explanations on commonsense QA, and found its high correlations with predicted answer accuracy, which is the opposite of token-level explainability that tends to get worse when the accuracy improves.
@@ 5 Discussion and Concluding Remarks
In our investigation of the token-level calibration of in-context learning in contemporary Large Language Models (LLMs), we delineate the intricate balance between ICL performance and calibration. Our findings underscore the importance of being circumspect in model deployment, as maximizing ICL performance does not invariably translate to improved calibration. As LMs continue to evolve and gain more capabilities such as having long enough context windows that can include the whole training set as in-context examples for some downstream tasks, our result can be pedagogical when users would like to examine their uncertainty through prediction probabilities. Moreover, the work suggests the following future directions:
@@ Acknowledgment
We thank Yu Bai, David Childers, Jean-Stanislas Denain for valuable feedback. Kempner Institute computing resources enabled this work.